OpenStudy (anonymous):

Please help with Trig Identities... tried it, but continued to get stuck at the end... Prove by working one side: (cos x/ 1+ sin x) - ( 1-sin x/ cos x)= 0

OpenStudy (amistre64):

cos(x)/(1+sin(x)) ?? is that corrrect?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

id assume yes, since cos(x)/1 would be futile :)

OpenStudy (amistre64):

add the right "term" to the right side of the (=) for starters

OpenStudy (amistre64):

cos x 1-sin x -------- = ------- 1+sin cos

OpenStudy (amistre64):

the RHS = sec - tan

OpenStudy (anonymous):

not = it is -...... it equals 0

OpenStudy (amistre64):

you should use algebraic techniques to "prove" the statement. All I did was move -((1-sin)/cos) to the other side for the moments :)

OpenStudy (amistre64):

if we cross multiply we get: cos^2 = (1-sin)(1+sin)

OpenStudy (anonymous):

i dont understand... im so confused

OpenStudy (amistre64):

cos^2 = 1 - sin^2 which is true

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

are you trying to just work one side and not manipulate it to your whim?

OpenStudy (anonymous):

yes only working one side to prove the other

OpenStudy (amistre64):

you could try decomposing fractions..... maybe if you want to try that way

OpenStudy (anonymous):

i ended up at: (1- sin ^2 x - 1- sin x)/ cos x

OpenStudy (amistre64):

cos 1-sin -------- - ------- = 0 1+sin cos ok....yeah, get like denominators and such...same thing I did pretty much: cos^2 - (1-sin^2) --------------- = 0 cos(1+sin)

OpenStudy (amistre64):

now we see that cos^2 - (cos^2) = 0 so that whole thing is 0

OpenStudy (amistre64):

do you see it?

OpenStudy (anonymous):

yes, thank you

OpenStudy (amistre64):

how did you get to this? (1- sin ^2 x - 1- sin x)/ cos x

OpenStudy (anonymous):

because cos ^2 = 1 - sin ^2

OpenStudy (anonymous):

thank you for helping me

OpenStudy (amistre64):

youre welcome :)